10. Calculate [H], [OH], pH for a 0.015 M aqueous solution of HCI. 11. Calculate [H], [OH], pH for a 0.45 M aqueous solution of NaOH. ヘ
Answer #1

10)

[H+] = [HCl] = 0.015 M

use:

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(0.015)

[OH-] = 6.67*10^-13 M

use:

pH = -log [H+]

= -log (0.015)

= 1.82

Answer:

[H+] = 0.015 M

[OH-]= 6.67*10^-13 M

pH = 1.82

I am allowed to answer only 1 question at a time